OpenStudy (anonymous):

Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each?

OpenStudy (anonymous):

I give medals =}

OpenStudy (anonymous):

Tony is 16, and Belinda is 40

OpenStudy (anonymous):

can you show me how you got that answer please?

OpenStudy (anonymous):

I just did trial and error - 16 x 2 = 32, + 8 =40. 40+16=56

OpenStudy (radar):

Tony and Belinda ages combined is 56, that means their sum of their ages is 56. If x is Tony's age then 56-x is Belindas age.

OpenStudy (anonymous):

I knew that part :) ^^

OpenStudy (radar):

It states that belinda age is twice tonys plus 8, now write an equation that will allow you to solve their ages.

OpenStudy (radar):

56-x=2x+8 that says Belinda's ages is twice tonys plus 8.

OpenStudy (radar):

3x=48 x=16 tonys age 40 is Belinda's